Solve: Helium gas has a volume of 250 mL at 0°C at 1.0 atm. What will be the final

pressure if the volume is reduced to 100 mL at 45°C?

Given: V1 = 250mL                                                       V2 = 100 mL

T1 = 00C + 273.15= 273.15 K                     T2 = 450C + 273.15=318.15 K P1 = 1 atm                                                                    P2 = ?


1
Expert's answer
2022-05-03T07:42:07-0400

Given:

P1 = 1.0 atm

V1 = 250 mL

T1 = 0°C + 273.15 = 273.15 K

P2 = ???

V2 = 100 mL

T2 = 45°C + 273.15 = 318.15 K


Solution:

Combined gas law can be used.

Mathematical expression for the combined gas law:

P1V1/T1 = P2V2/T2

Cross-multiply to clear the fractions:

P1V1T2 = P2V2T1

Divide to isolate P2:

P2 = P1V1T2 / V2T1

P2 = (1.0 atm × 250 mL × 318.15 K) / (100 mL × 273.15 K) = 2.9 atm

P2 = 2.9 atm


Answer: The final pressure will be 2.9 atm